OpenStudy (anonymous):

what is x-2y+3z if x=-2,y=-3 and z=1

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

x-2y+3z gives: -2 - 2*(-3) + 3*1 = -2 -(-6) + 3 = -2 + 6 + 3 = 7

OpenStudy (anonymous):

thats what i got but the answer key is saying -12

OpenStudy (anonymous):

oh my bad. I didn't see the -3

OpenStudy (anonymous):

That's a very quite answer in my opinion. Are there maybe any typo's you made?

OpenStudy (anonymous):

no typos, i'm sticking with 7 and assume the answer key is wrong. thanks bro